# A Multi-Task Foundation Model for Wireless Channel Representation Using Contrastive and Masked Autoencoder Learning

### v1

@@ III Proposed Architecture
Figure 1: Illustration of the WiMAE and ContraWiMAE architectures, detailing preprocessing steps, encoder processing, and decoding. WiMAE does not include the contrastive head block with dashed lines.
In this section, we detail our proposed WiMAE and ContraWiMAE architectures to model the encoder-decoder pair $(f_{\theta},g_{\phi})$ and the contrastive operator $\mathcal{G}_{\psi}$. The main building blocks are illustrated in Fig. 1, in which the contrastive head block, pictured with dashed lines, is the only difference between WiMAE and ContraWiMAE.
@@ III-A Preprocessing Operations
Next, we transform the visible patches into a sequence representation by ordering them based on their original positions in a row-major order. Let $\mathbf{p}_{\text{v}}^{k}$ be the $k$-th visible patch in this ordering, with $1\leq k\leq N_{\text{v}}$, where $N_{\text{v}}=(1-m_{\text{r}})\cdot K$ is the number of visible patches. To handle the complex-valued nature of the channel matrix, we separate each patch into its real and imaginary components $(\mathbf{p}_{\text{v}}^{k})^{\text{r}}=\Re(\mathbf{p}_{\text{v}}^{k})$ and $(\mathbf{p}_{\text{v}}^{k})^{\text{i}}=\Im(\mathbf{p}_{\text{v}}^{k})$, respectively, each with dimension $N_{\text{p,s}}\times N_{\text{p,f}}$. We flatten each component to a vector of length $d_{\text{p}}$ and construct the visible patch sequence:
@@
To incorporate positional information, we maintain a set of positional encodings $\mathbf{PE}_{\text{e}}\in\mathbb{R}^{2K\times d_{\text{e}}}$ for all possible patch positions. The first $K$ rows correspond to the positions of real components while the last $K$ rows correspond to the positions of imaginary components. Let $\mathcal{I}_{\text{v}}\subset\{1,2,...,2K\}$ be the set of indices in the original patch sequence corresponding to visible patches (both real and imaginary components). We then add the subset of positional encodings corresponding to these visible patches as follows:
@@ III-B Encoder Architecture
The outputs from all $M_{\text{enc}}$ attention heads are concatenated and linearly projected to obtain the final MHSA output:
@@
For each layer $l$, the output $\mathbf{Z}^{(l)}=\mathbf{Z}_{\text{out}}$ becomes the input to the next layer: $\mathbf{Z}_{\text{in}}=\mathbf{Z}^{(l)}$ for layer $l+1$. After passing through all transformer layers, the final output $\mathbf{Z}^{(L_{\text{enc}})}$ becomes the encoder output $\mathbf{Z}_{\text{enc}}\in\mathbb{R}^{2N_{\text{v}}\times d_{\text{e}}}$ that captures the contextual information from visible patches while maintaining their positional relationships.
@@ III-C Decoder Architecture
III-D ContraWiMAE Extension
Building upon the WiMAE architecture described above, we implement a contrastive learning extension that operates in parallel with the reconstruction pathway. This section details the specific implementation aspects of the contrastive component.
@@ IV Methodology
This section details the methodology used in our experimental evaluation of WiMAE and ContraWiMAE, including dataset generation, pretraining setup, and adaptations for downstream tasks. Through these experiments, we aim to demonstrate the effectiveness of our self-supervised learning approaches for wireless channel representation learning. A thorough discussion of the results is provided in Section V.
@@ IV-A Dataset Generation
Using the DeepMIMO dataset [41], we generate 1.14M samples at 3.5 GHz operating frequency from 15 scenarios for pretraining. Each sample in the dataset corresponds to a channel for a distinct user and a single base station from one of the scenarios from: Seattle, Los Angeles, Chicago, Houston, Phoenix, Philadelphia, Miami, Dallas, San Francisco, Austin, Columbus, New York, ASU Campus, Boston5G and O1. We also generate a downstream task dataset with 14,840 samples from 6 unseen scenarios: Denver, Fort Worth, Oklahoma, Indianapolis, Santa Clara, and San Diego. Similar to the pretraining dataset, each scenario features one base station with uniformly distributed users. For a fair comparison, we match the configuration in both the pretraining and downstream datasets to the one in LWM [33].
IV-B Pretraining of WiMAE
We train WiMAE for 3,000 epochs with an 80/20 training/validation split on the pretraining dataset. We use the Adam optimizer with batch size of 3,072 and learning rate decaying from $3\times 10^{-4}$ to $3\times 10^{-6}$ with cosine decay. We explore different masking ratios $m_{\text{r}}\in\{0.45,0.60,0.75,0.90\}$, network depths $L_{\text{enc}}\in\{6,12,18\}$ and $L_{\text{dec}}\in\{4,8,12\}$, encoding dimensions $d_{\text{e}}\in\{32,64,96,128\}$, and patch configurations $(N_{\text{p,s}},N_{\text{p,f}})\in\{(1,16),(4,4)\}$, with fixed attention heads $M_{\text{enc}}=16$ and $M_{\text{dec}}=8$, as shown in Fig. 2.
TABLE I: WiMAE and ContraWiMAE pretraining configurations.
Common parameters Value
Training/validation split 80/20
Epochs 3,000
Batch size 3,072
Optimizer Adam
Learning rate $3\times 10^{-4}$ with cosine decay
Attention heads $M_{\text{enc}}=16$, $M_{\text{dec}}=8$
WiMAE Value
Masking ratios $m_{\text{r}}$0.45, 0.60, 0.75, $\{0.45,\ \mathbf{0.60},\ 0.75,\ 0.90\}$
Encoder depth $L_{\text{enc}}$6, 12, $\{6,\ \mathbf{12},\ 18\}$
[…]
Embedding dim. $d_{\text{e}}$32, 64, 96, $\{32,\ \mathbf{64},\ 96,\ 128\}$
Patch size $(N_{\text{p,s}},N_{\text{p,f}})$$\{\mathbf{(1,16)},\ (4,4)\}$
ContraWiMAE Value
Initialization Warm-start from WiMAE
Loss coefficient in (4) $\alpha=0.9$
Contrastive temperature in (3) $\tau=0.1$
Contrastive dimension $d_{\text{c}}$ 64
Positive pair generation AWGN to achieve SNR $\in$ [20, 30] dB
Note: Boldface values are chosen for downstream task evaluation.
(a) Top-1 accuracy vs. mask ratio.
(b) Top-1 accuracy vs. model depth.
(c) Top-1 accuracy vs. dimension.
Figure 2: Beam prediction performance for WiMAE with linear probing across different model parameters and codebook sizes.
The best reconstruction performance, illustrated in Fig. 3, is obtained with $(N_{\text{p,s}},N_{\text{p,f}})=(4,4)$, achieving 29.14 dB SNR on the validation set. However, $(N_{\text{p,s}},N_{\text{p,f}})=(1,16)$ leads to better performance in downstream tasks and is considered the default for the rest of the paper. Table I summarizes the pretraining configurations used for WiMAE and ContraWiMAE.
Figure 3: Magnitude (left) and phase (right) reconstruction for WiMAE, obtained with patch size $(N_{\text{p,s}},N_{\text{p,f}})=(4,4)$. Each row is a channel sample. Columns correspond to original, masked, reconstructed, and masked reconstructed channels.
@@ IV-B Pretraining of WiMAE
For comparisons with baseline models on downstream tasks, we set $m_{\text{r}}=0.60$, $L_{\text{enc}}=12$, $L_{\text{dec}}=4$, $d_{\text{e}}=64$, and $(N_{\text{p,s}},N_{\text{p,f}})=(1,16)$. These choices result in 31% fewer encoder parameters and 7% fewer total parameters compared with LWM (see Table II). For LWM, we use publicly available pretrained weights and the code to generate its embeddings [33] without retraining or any modifications.
@@ IV-C Pretraining of ContraWiMAE
We warm-start the encoder and decoder components of ContraWiMAE by copying the corresponding weights from WiMAE to preserve its reconstruction capability. We use the same learning rate schedule, optimizer, and batch size for continual pretraining of ContraWiMAE on the same pretraining dataset. To maximize the transferability of the reconstruction capability, we use $m_{\text{r}}=0.60$, $L_{\text{enc}}=12$, $L_{\text{dec}}=4$, $d_{\text{e}}=64$, and $(N_{\text{p,s}},N_{\text{p,f}})=(1,16)$, matching the parameters for WiMAE used in downstream task evaluations. We set the reconstructive and contrastive loss weights to $\alpha=0.9$ and $0.11-\alpha=$ 0.1, respectively, using a temperature parameter $\tau=0.1$. We also set the contrastive space dimension to $d_{\text{c}}=64$ and uniformly sample the noise variance for positive pairs to achieve channel SNR values between 20 dB and 30 dB.
TABLE II: Parameter counts of different configurations.
@@
Note: CWiMAE = ContraWiMAE (contrastive learning variant). Model names show ($L_{\text{enc}}$,$L_{\text{dec}}$) where $L_{\text{enc}}$ and $L_{\text{dec}}$ are encoder and decoder layers. For WiMAE and CWiMAE: $M_{\text{enc}}{=}16$, $M_{\text{dec}}{=}8$, $d_{\text{e}}{=}64$.
@@ IV-D Downstream Task Adaptation
We evaluate WiMAE and ContraWiMAE on two downstream tasks: (i) cross-frequency beam selection and (ii) line-of-sight (LoS) detection, using 70/20/10 training/validation/test splits. The beam selection task evaluates how well each encoding captures the channel’s detailed site-specific structure across antenna elements and subcarriers, while the LoS task measures the effectiveness of the extracted features as compact channel summaries.
@@ IV-D1 Cross-frequency Beam Selection
The beam selection task predicts optimal mmWave beams at 28 GHz from 3.5 GHz channel data using predefined codebook sizes (CS) of 16, 32, 64, 128, and 256. This reduces the channel estimation overhead that is normally needed in beam management [42] while testing how well the learned representations capture cross-frequency propagation characteristics. We report top-1 and top-3 accuracy: top-1 accuracy measures the proportion of test samples where the predicted best beam matches the true optimal beam; top-3 accuracy extends this by considering a prediction correct if the true optimal beam is among the top three highest-scoring predicted beams. We compare models trained on WiMAE and ContraWiMAE embeddings with models trained on LWM embeddings and models trained on raw channel data without any representation learning (RAW), as summarized in Tables III and IV.

### v2

@@ III Proposed Architecture
Figure 1: Illustration of the ContraWiMAE blocks, detailing preprocessing steps, encoder-decoder processing, and the contrastive head.
In this section, we detail our proposed architecture to model the encoder-decoder pair $(f_{\theta},g_{\phi})$ and the contrastive operator $\mathcal{G}_{\psi}$. The main building blocks are illustrated in Fig. 1, in which the contrastive head block, pictured with dashed lines, is the only difference between reconstruction-only baseline and ContraWiMAE.
@@ III-A Preprocessing Operations
Next, we transform the visible patches into a sequence representation by ordering them based on their original positions in a row-major order. Let $\mathbf{p}_{\text{v}}^{k}$ be the $k$-th visible patch in this ordering, with $1\leq k\leq N_{\text{v}}$, where $N_{\text{v}}=(1-m_{\text{r}})\cdot K$ is the number of visible patches. To handle the complex-valued nature of the channel matrix, we separate each patch into its real and imaginary components rather than magnitude and phase to preserve linearity and avoid discontinuities at phase wrapping boundaries. We define $(\mathbf{p}_{\text{v}}^{k})^{\text{r}}=\Re(\mathbf{p}_{\text{v}}^{k})$ and $(\mathbf{p}_{\text{v}}^{k})^{\text{i}}=\Im(\mathbf{p}_{\text{v}}^{k})$, respectively, each with dimension $N_{\text{p,s}}\times N_{\text{p,f}}$. We flatten each component to a vector of length $d_{\text{p}}$ and construct the visible patch sequence:
@@
To incorporate positional information, we maintain a set of positional encodings $\mathbf{PE}_{\text{e}}\in\mathbb{R}^{2K\times d_{\text{e}}}$ for all possible patch positions. The first $K$ rows correspond to the positions of real components while the last $K$ rows correspond to the positions of imaginary components. Let $\mathcal{I}_{\text{v}}\subset\{1,2,...,2K\}$ be the set of indices in the original patch sequence corresponding to visible patches (both real and imaginary components). We add the subset of positional encodings corresponding to these visible patches as follows:
@@ III-B Encoder Architecture
The outputs of all $M_{\text{enc}}$ attention heads are concatenated and linearly projected to obtain the final MHSA output:
@@
For each layer $l$, the output $\mathbf{Z}^{(l)}=\mathbf{Z}_{\text{out}}$ becomes the input to the next layer: $\mathbf{Z}_{\text{in}}=\mathbf{Z}^{(l)}$ for layer $l+1$. After passing through all transformer layers, the final output $\mathbf{Z}^{(L_{\text{enc}})}$ becomes the encoder output $\mathbf{Z}_{\text{enc}}\in\mathbb{R}^{2N_{\text{v}}\times d_{\text{e}}}$ that captures contextual information from visible patches while maintaining their positional relationships.
@@ III-C Decoder Architecture
III-D Contrastive Head
Contrastive head operates in parallel with the reconstruction pathway. This section details the specific implementation aspects of the contrastive component.
@@ IV Methodology
This section details the methodology used in our experimental evaluation of ContraWiMAE, including dataset generation, pretraining setup, and adaptations for downstream tasks. Through these experiments, we aim to demonstrate the effectiveness of our self-supervised learning approach for wireless channel representation learning. A thorough discussion of the results is provided in Section VI.
@@ IV-A Dataset Generation
For our experimental evaluation, we configure the system with $N_{\text{f}}=32$ subcarriers, $\Delta f=30$ kHz subcarrier spacing (total bandwidth $B=960$ kHz), and $N_{\text{s}}=32$ antennas with half-wavelength spacing. All antennas have isotropic radiation patterns and single polarization, with a maximum of 20 propagation paths considered. Using the DeepMIMO dataset [42], we generate 2.5M samples at 3.5 GHz operating frequency from 56 scenarios for pretraining. Each sample in the dataset corresponds to a channel for a distinct user-base station pair from one of the following scenarios: Amsterdam, ASU Campus, Athens, Bangkok, Barcelona, Beijing, Boston, Brussels, Cairo, Cape Town, Charlotte, Chicago, Denver, Dubai, Edinburgh, Florence, Fort Worth, Fujiyoshida, Granada, Hatsukaichi, Helsinki, Hong Kong, Indianapolis, Istanbul, Jerusalem, Kyoto, Havana, Lisbon, Los Angeles, Madrid, Marrakesh, Mumbai, New Delhi, New York, North Jakarta, Oklahoma City, Philadelphia, Phoenix, Reykjavik, Rio de Janeiro, Rome, San Francisco, San Nicolas, Saint Petersburg, Santa Clara, Santiago, Seattle, Seoul, Singapore, Stockholm, Sumida City, Sydney, Taipei, Taito City, Toronto, and Warsaw. We also generate a dataset for downstream tasks with 0.55M total samples from 10 unseen scenarios: Austin, Centro, Columbus, Dallas, Gurbchen, Houston, Miami, Montreal, Prague, and San Diego. Similar to the pretraining dataset, some scenarios feature more than one base station with uniformly distributed users. Apart from the 3.5 GHz channels, we generate channels at 28 GHz for downstream task scenarios with available ray-tracing data. To ensure fairness in comparisons involving LWM, we also use a smaller set of scenarios exactly matching the configuration described in the respective work [34] for both the pretraining and downstream datasets. This dataset consists of 1.14M samples for pretraining and 14,840 samples for downstream tasks.
We selected the DeepMIMO dataset for three key reasons: (i) to enable direct benchmarking with prior work such as LWM [34] under identical input features and data-generation pipelines, (ii) to ensure reproducibility by providing a publicly available dataset that enables other researchers to retrain and validate our approach, and (iii) to test generalization across diverse geographical layouts, leveraging DeepMIMO’s extensive collection of over 50 ray-traced scenarios spanning cities worldwide.
IV-B Reconstruction-only Pretraining
For brevity, we refer to our reconstruction-only foundation model as WiMAE in the following sections. We train WiMAE for 3,000 epochs with a 90/10 training/validation split and a batch size of 8,096. We use the AdamW optimizer with linear rate warm-up reaching $3\times 10^{-4}$ in the first 10 epochs, followed by cosine decay to $3\times 10^{-6}$.
@@ IV-B Reconstruction-only Pretraining
While we explore many different configurations, as discussed in detail in Section VI, for the comparisons with baseline models on the downstream tasks, we set $m_{\text{r}}={\color[rgb]{0,0,0}0.90}$, $L_{\text{enc}}=12$, $L_{\text{dec}}=4$, $d_{\text{e}}=64$, and $(N_{\text{p,s}},N_{\text{p,f}})={\color[rgb]{0,0,0}(16,1)}$, striking a balance between complexity and performance. These choices result in 31% fewer encoder parameters and 7% fewer total parameters compared with LWM (see Table I). Here, we note that the choice of $m_{\text{r}}$ does not affect the number of parameters, and hence the model expressivity. However, it directly affects the pretext task difficulty, necessitating searching for the optimal mask ratio for each patch shape for optimal learning. For LWM, we use publicly available pretrained weights and the code to generate its embeddings [34] without retraining or modifications.
@@ IV-C Pretraining of ContraWiMAE
We warm-start the encoder and decoder components of ContraWiMAE from WiMAE. We use the same learning rate schedule, optimizer, and batch size for continual pretraining of ContraWiMAE on the same pretraining dataset. To maximize the transferability of the reconstruction capability, we use the same parameters as those used in the reconstruction foundation.
We set the reconstructive-contrastive trade-off weight $\alpha=0.9$ and the temperature parameter $\tau=0.2$. We also set the contrastive space dimension to $d_{\text{c}}=64$ and uniformly sample the noise variance for positive pairs to achieve channel SNR values between 5 dB and 40 dB. We choose these contrastive learning parameters through ablation studies based on the total loss in Eq. (4).
TABLE I: Parameter counts of different configurations.
@@
Note: CWiMAE = ContraWiMAE. Model names show ($L_{\text{enc}}$$L_{\text{dec}}$) where $L_{\text{enc}}$ and $L_{\text{dec}}$ are encoder and decoder layers. For WiMAE and CWiMAE: $M_{\text{enc}}{=}16$, $M_{\text{dec}}{=}8$, $d_{\text{e}}{=}64$.
@@ IV-D Downstream Task Adaptation
We evaluate ContraWiMAE on three downstream tasks: (i) cross-frequency beam selection, (ii) line-of-sight (LoS) detection, and (iii) channel estimation. For LoS detection and beam selection, we perform both parametric and non-parametric evaluations where we freeze the pretrained encoder. For parametric evaluation, which adds expressivity to the prediction pipeline, we train a linear classifier on top of the encoder representations, following standard practices. We also train custom residual networks to see the marginal improvement possible through learning task-specific non-linearities. For our non-parametric evaluation, we assess the performance of a kNN classifier on the pretrained encoder representations, disentangling representation quality from task head capacity. For channel estimation, we keep the encoder-decoder pair and measure the training-free performance. We also analyze the gain from fine-tuning the model on this specific task.
TABLE II: Top-1/top-3 accuracy for beam selection with linear probing.
Model Codebook Size (CS) Training Budget
1% 2% 5% 10% 25% 50% 100%
WiMAE 16 22.622.6/$36.1$ 32.232.2/$48.5$ 2727/$42.7$ 31.731.7/$43.5$ 36.136.1/$50.2$ 38.938.9/$49.5$ 41.441.4/$51$
32 15.615.6/$28.3$ 18.918.9/$26.4$ 19.819.8/$28$ 23.823.8/$32.8$ 28.228.2/$34$ 3333/$37.2$ 31.531.5/$36.6$
64 1313/$19.5$ 14.414.4/$23.1$ 1313/$18.3$ 17.817.8/$22.6$ 24.124.1/$29.8$ 26.326.3/$31.7$ 26.626.6/$30.4$
128 7.27.2/$10.3$ 9.69.6/$15.1$ 10.210.2/$15.7$ 14.214.2/$18.7$ 18.918.9/$22$ 19.119.1/$23.4$ 16.616.6/$19.1$
256 55/$8.4$ 6.76.7/$11.5$ 6.96.9/$11.2$ 9.49.4/$13.7$ 10.510.5/$13.7$ 12.312.3/$14.8$ 10.210.2/$11.8$
ContraWiMAE 16 49.2/69.7 55.9/75.7 62.7/79.4 72.1/87.2 72.2/83.6 77.4/87.5 79/89.5
32 35.9/47.4 41.4/57.1 50.6/68.2 63.5/77.5 68.4/76.3 68.3/76.6 75.3/84.2
64 24.7/36.2 29/44.6 31.3/46.3 40.1/54.3 48.3/62.8 56.1/69.7 64.5/79
128 14/20.6 16.2/27.7 18.5/30.3 23.7/34.7 29/39.5 35.4/46.3 39.3/49.6
256 8/14.5 10.9/18.7 11.811.8/21.5 15.2/$23.3$ 18/$26.6$ 18.9/$26.1$ 20.220.2/$26.6$
LWM 16 16.316.3/$31.5$ 22.422.4/$33.7$ 18.418.4/$33$ 25.325.3/$36.1$ 36.736.7/$46.2$ 3737/$47$ 44.144.1/$54.6$
32 16.416.4/$25.3$ 20.120.1/$27.2$ 14.214.2/$22.6$ 25.725.7/$34.9$ 3636/$44.3$ 37.137.1/$45.1$ 39.639.6/$46.8$
64 14.814.8/2121 18.518.5/$25.9$ 12.112.1/$17.5$ 2020/$26.6$ 26.226.2/$33.3$ 32.432.4/$39$ 35.435.4/$41.5$
128 99/$14$ 10.710.7/$18.4$ 7.67.6/$12.7$ 12.312.3/$19.5$ 17.517.5/$23.3$ 21.621.6/$27.8$ 23.723.7/$29.4$
256 66/$9.5$ 77/$11.8$ 5.35.3/$9.1$ 7.67.6/$11.7$ 11.111.1/$16.2$ 12.412.4/$18$ 12.512.5/$17.7$
RAW 16 31.731.7/$40.3$ 31.731.7/$41.8$ 30.830.8/$41.6$ 38.138.1/$46.2$ 39.139.1/$47.1$ 38.938.9/$44.9$ 41.241.2/$47.4$
32 27.627.6/$38.5$ 24.524.5/$36.7$ 34.134.1/$43.3$ 36.836.8/$46.4$ 36.536.5/$43.8$ 41.241.2/$47$ 42.542.5/$48.1$
64 19.419.4/$31.1$ 22.622.6/$34.3$ 27.627.6/$40.8$ 30.530.5/$42.4$ 33.533.5/$43.5$ 33.633.6/$42.6$ 36.136.1/$43.7$
128 9.99.9/$18.2$ 12.712.7/$23.4$ 17.417.4/$29.5$ 19.919.9/$31.8$ 23.923.9/$35$ 2727/$38.5$ 29.329.3/$37.8$
256 5.95.9/$11.7$ 99/$16$ 12/$20.1$ 13.513.5/23.8 16.616.6/29 17.117.1/29.6 20.9/31.6
Note: RAW refers to models trained directly on raw channel data without any representation learning. Boldface indicates the best result for each (CS, training budget) pair.
These tasks probe complementary aspects of the learned representations. The beam selection task evaluates how well each encoding captures the channel’s detailed site-specific structure across antenna elements and subcarriers, while the LoS detection measures the effectiveness of the extracted features as compact channel summaries. Channel estimation assesses reconstruction capability by evaluating the model’s ability to recover the underlying channel from sparse and noisy observations.
@@ IV-D1 Cross-frequency Beam Selection
The beam selection task predicts optimal mmWave beams at 28 GHz from 3.5 GHz channel data using predefined codebooks with uniform beam patterns and codebook sizes (CS) of 16, 32, 64, 128, and 256. This reduces the channel estimation overhead that is normally needed in beam management [43] while testing how well the learned representations capture cross-frequency propagation characteristics. We report top-1 and top-3 accuracy: top-1 accuracy measures the proportion of test samples for which the predicted best beam matches the true optimal beam (the beam that maximizes the received power); top-3 accuracy extends this by considering a prediction correct if the true optimal beam is among the top three highest-scoring predicted beams. We compare models trained on ContraWiMAE embeddings with models trained on LWM, WiMAE, and supervised embeddings and models trained on raw channel data without any representation learning (RAW), as summarized in Tables II and III and several figures in Section VI. The supervised baseline consists of an encoder (same as ContraWiMAE) and a linear classifier attached to it. However, we train this model directly on the pretraining dataset with 3.5 GHz channel-beam pairs separately for each codebook size. We discard the linear classifier and use the encoder representations for downstream tasks.